4.LVMH is a French firm listed on the Paris Bourse, You can access the central limit order book directlv on the internet and find the following information (the limit Prices for sell orders are ask prices and those for buy orders are bid prices: Sell Orders Buy Orders Quantity LimiQuantity Limit 1,000 58 3,000 54 1,000 53 1,000 52 2,000 46 500 51 10049 500 48 1,000 47 10,000 44 You wish to buy 3,000 shares and enter a market order to buv those shares. A Market order will be executed against the best matching order, At what Price will vou buy the shares? If you wish to sell 500 shares, at what price will you sell the shares?
Answer #1

Solution:

Part A ) Entering market order to buy 3,000 shares: When one enters the market order then the order will be executed starting from the lowest offered price.

In the given question lowest offer price is 51 and next price is 52 and so on.

Price: 51 Quantity: 500

Price: 52 Quantity: 1000

Price: 53 Quantity: 1000

Price: 54 Quantity: 500

Average Price =( 51 * 500 + 52 *1000 + 53*1000 + 54*500 ) / 3000 = 157500 /3000 = 52.5

Part B) Selling 500 shares at market order: When market order to sell is placed then the highest bid price order will be fetched first.

Here highest bid price is 49 and quantity is 100.

Next highest bid is 48 and quantity is 500. But in order to fulfill the market order only 400 quantity is needed

Executed order will be

Price : 49 Quantity: 100

Price: 48 Quantity: 400

Average price = (49*100 + 48*400)/500 = 24100/500 = 48.2